# Adaptive Refinement Protocols for Distributed Distribution Estimation under $\ell^{p}$-Losses

### v1

@@ Abstract
Consider the communication-constrained estimation of discrete distributions under $\ell^{p}$ losses, where each distributed terminal holds multiple independent samples and uses limited number of bits to describe the samples. We obtain the minimax optimal rates of the problem in most parameter regimes. An elbow effect of the optimal rates at $p=2$ is clearly identified. To show the optimal rates, we first design estimation protocols to achieve them. The key ingredient of these protocols is to introduce adaptive refinement mechanisms, which first generate rough estimate by partial information and then establish refined estimate in subsequent steps guided by the rough estimate. The protocols leverage successive refinement, sample compression and thresholding methods to achieve the optimal rates in different parameter regimes. The optimality of the protocols is shown by deriving compatible minimax lower bounds.
@@ I Introduction
  • For $p>2$, we introduce additional sample compression methods to aid the adaptive refinement procedure. The methods compress the description for samples and reduce the communication budget, allowing more samples to be transmitted within limited budget. The resulting protocols can achieve the optimal rates for both relatively large $n$ and $n=1$. To show the optimality for $n=1$, we further establish a compatible lower bound that is strictly better than that in [15, 16],
@@ III-D Optimal Rates for $p>2$ and $n=1$
For $n=1$ and $p>2$, the lower bound can be derived by specializing Lemma 2, and the compatible (up to logarithmic factors) upper bound is shown in the following thoerem.
@@
Let $p>2$ and $n=1$. If $m(2^{l}\wedge k^{\frac{2}{p}})\geq k^{2}$, then $R(m,1,k,l,p)\preceq\frac{k\log^{\frac{p}{2}}k}{(m2^{l})^{\frac{p}{2}}}\vee\frac{1}{m^{\frac{p}{2}}}$ and $R(m,1,k,l,p)\succeq\frac{k}{(m2^{l})^{\frac{p}{2}}}\vee\frac{1}{m^{\frac{p}{2}}}$.
@@
The upper bound is by Proposition 5 in Section VIII. We sketch the proof here. The difficulty is mainly in the construction of sample compression methods for the refinement step of the protocol. Entries $p_{W}(w)>\frac{1}{k}$ are divided into $\log k$ groups based on their order. Entries in the same group are then estimated simultaneously by invoking the simulation protocol in [5]. Through such grouping and compression methods, more samples can be simulated for relatively larger entries $p_{W}(w)$. Hence the optimal rates (up to logarithmic factors) is obtained. See Section VIII for details of the protocol and its analysis. ∎
@@
Note that the central bound $\frac{1}{m^{\frac{p}{2}}}$ without the communication constraints is neglected by previous works [15, 16] (see Theorem 6 in [15] and Corollary 3.2 in [16]). Hence the lower bounds in both works are clearly not tight (for $p>2$). The work [16] further claimed that the lower bound $\frac{k}{(m2^{l})^{\frac{p}{2}}}\vee\frac{k^{1-\frac{p}{2}}}{m^{\frac{p}{2}}}$ is optimal (see Lemma 3.3 therein), but the sketch given there is too brief and not sufficient to describe a protocol that achieves the bound. In fact, given that the lower bound in [16] can be strictly improved, it is impossible to show its optimality. Moreover, constructing the protocol that achieves the optimal rates for $p>2$ is not that straightforward and needs additional ideas.
@@ VIII The Protocol for $n=1$
In this section, we design an adaptive refinement protocol with sample compression that achieves the optimal rate for $n=1$. Similar to the discussion in Remark 5, it suffices to show the following proposition for $p\geq 2$.
@@
Let $p\geq 2$, $n=1$, $m2^{l}\geq k^{2}$ and $2^{l}\leq k^{\frac{2}{p}}$. Then for the problem in Section II, there exists an interactive protocol such that for any $\bm{p}_{W}\in\Delta_{\mathcal{W}}$, the protocol outputs an estimate $\hat{\bm{p}}_{W}$ satisfying $\mathbb{E}[\|\hat{\bm{p}}_{W}-\bm{p}_{W}\|_{p}^{p}]=O\left(\frac{k\log^{\frac{p}{2}}(\frac{k}{2^{l}})}{(m2^{l})^{\frac{p}{2}}}\right)$.
We want to exploit the simulation protocol in [19]. With a probability of at least $\frac{3}{4}$, the protocol simulates one sample from a $k$-dimensional distribution using $O(\frac{4k}{2^{l}})$ encoders, where each of them is restricted to send a message of $l$-bits. Then all the encoders can be divided into groups and each group can cooperatively simulate one sample. Hence the following lemma can be obtained by the Hoeffding’s inequality.
Lemma 6.
Suppose that there are $m^{\prime}$ users and each of them observe an i.i.d. sample from a $k^{\prime}$-dimensional distribution. If $2^{l}\leq k^{\prime}$, then there exists a randomized protocol $\mathrm{DistrSim}(m^{\prime},k^{\prime},l)$ that generates $\frac{m^{\prime}2^{l}}{8k^{\prime}}$ i.i.d. random samples at the decoder side with probability at least $1-e^{-\frac{m^{\prime}2^{l}}{32k^{\prime}}}$.
@@ VIII-A Motivation of the Protocol
The most natural idea is to first invoke the simulation protocol to output $M=O(\frac{m2^{l}}{k})$ samples from the distribution $\bm{p}_{\mathcal{W}}$ at decoder side; then estimate $\bm{p}_{\mathcal{W}}$ using $M$ samples by a traditional central estimation method. It can achieve the optimal minimax rate $\frac{k}{m2^{l}}$ for $p=2$, and hence the optimal rate $\frac{k}{(m2^{l})^{\frac{p}{2}}}$ for $1\leq p\leq 2$. However, for $p\geq 2$, using $M$ i.i.d. samples to estimate the underlying distribution under the $\ell^{p}$ loss can only achieve a rate of $\frac{1}{M^{\frac{p}{2}}}=(\frac{k}{m2^{l}})^{\frac{p}{2}}$ , which leaves a gap with the lower bound $\frac{k}{(m2^{l})^{\frac{p}{2}}}$ by Lemma 1. The above naive protocol is not optimal and we can show that the lower bound $\frac{k}{(m2^{l})^{\frac{p}{2}}}$ is optimal.
@@
An observation is that, given $M$ samples from a distribution $\bm{p}_{\mathcal{W}}$, the estimation error for each entry $p_{\mathcal{W}}(w)$ is of order $\left(\frac{p_{W}(w)(1-p_{W}(w))}{M}\right)^{\frac{p}{2}}$, which depends on $p_{\mathcal{W}}(w)$. For larger $p_{\mathcal{W}}(w)$, the number $M$ of samples should be larger. So the simulation protocol should be repeated more for the estimation of such $p_{\mathcal{W}}(w)$, while less for smaller $p_{\mathcal{W}}(w)$. The intuition can be achieved by the following two-step estimation procedure. In the first step, a rough estimate is obtained. The estimation is used to direct the resource allocation in the next step. In the second step, a sample compression method is designed to save the communication budget, so that more samples are transmitted for larger $p_{\mathcal{W}}(w)$.
Details of the protocol are presented as follows, and the error analysis can be found in Appendix F.
VIII-B The Adaptive Refinement Protocol with Sample Compression for $n=1$
VIII-B1 Rough Estimation
Let the first $\frac{m}{2}$ encoders and the decoder invoke the protocol $\mathrm{DistrSim}(\frac{m}{2},k,l)$ to simulate the samples from the distribution $\bm{p}_{\mathcal{W}}$. The protocol generates $M_{0}\triangleq\frac{m2^{l}}{16k}$ i.i.d. random samples $(W^{0}_{l})_{l=1}^{M_{0}}$ at the decoder side, except in the case that an unsuccessful event $\mathcal{E}_{0}$ with probability $\mathbb{P}[\mathcal{E}_{0}]\leq e^{-\frac{m2^{l}}{64k}}$ occurs.
If the above simulation succeeds, then for each $w\in\mathcal{W}$, let
$\hat{p}_{W}^{1}(w)=\frac{\sum_{l=1}^{M_{0}}\mathds{1}_{W^{0}_{l}=w}}{M_{0}}.$
Otherwise, let $\hat{p}_{W}^{1}(w)=\frac{1}{k}$ for each $w\in\mathcal{W}$.
VIII-B2 Adaptive Refinement with Sample Compression
Divide all these encoders into $\lceil\log(\frac{k}{2^{l}})\rceil$ groups and each group has $m^{\prime}=\lfloor\frac{m}{2\lceil\log(\frac{k}{2^{l}})\rceil}\rfloor$ encoders. All the remaining $\frac{m}{2}$ encoders and the decoder compute
$\mathcal{W}^{1}=\left\{w\in\mathcal{W}:\hat{p}^{1}_{W}(w)\in\left[\frac{1}{2^{l+1}},1\right]\right\},$
$\mathcal{W}^{s}=\left\{w\in\mathcal{W}:\hat{p}^{1}_{W}(w)\in\left[\frac{1}{2^{l+s}},\frac{1}{2^{l+s-1}}\right)\right\},s=2,...,\lceil\log(\frac{k}{2^{l}})\rceil-1,$
$\mathcal{W}^{\lceil\log(\frac{k}{2^{l}})\rceil}=\mathcal{W}-\left(\cup_{s=1}^{\lceil\log(\frac{k}{2^{l}})\rceil-1}\mathcal{W}^{s}\right),$
where it is immediate that $|\mathcal{W}^{s}|\leq 2^{l+s}$.
For the $i$-th encoder in the $s$-th group, it computes $W^{s}_{i}=h_{s}(W_{i})$, where $W_{i}$ is its observed sample and
$h_{s}(w)=\left\{\begin{aligned} &w,w\in\mathcal{W}^{s},\\ &\emptyset,w\notin\mathcal{W}^{s}.\end{aligned}\right.$
Let $W^{s}=h_{s}(W)$ and $p_{W^{s}}$ be its distribution of dimension no more than $12^{l+s}+$ 1. Then each encoder in the $s$-th group holds one sample $W^{s}_{i}\sim p_{W^{s}}$. Let these encoders and the decoder invoke the protocol $\mathrm{DistrSim}(m^{\prime},|\mathcal{W}^{s}|+1,l)$ to simulate samples from the distribution $\bm{p}_{\mathcal{W}^{s}}$. The protocol generates $M_{s}\triangleq\frac{m^{\prime}2^{l}}{8(2^{l+s}+1)}$ i.i.d. random samples $(\tilde{W}^{s}_{l})_{l=1}^{M_{s}}$ at the decoder side, except when an unsuccessful event $\mathcal{E}_{s}$ with probability $\mathbb{P}[\mathcal{E}_{s}]\leq e^{-\frac{m^{\prime}2^{l}}{32(2^{l+s}+1)}}$ occurs.
If the above simulation succeeds, then for each $w\in\mathcal{W}^{s}$ and each $s$, let
$\hat{p}_{W}^{2}(w)=\frac{\sum_{l=1}^{M_{s}}\mathds{1}_{\tilde{W}^{s}_{l}=w}}{M_{s}}.$
Otherwise, let $\hat{p}_{W}^{2}(w)=\frac{1}{k}$ for each $w\in\mathcal{W}^{s}$ and each $s$. The decoder then outputs the refined estimate $\hat{\bm{p}}_{W}^{2}$.
@@ IX Lower Bounds
Lemma 7 is proved by exploiting the results for $p=1$ in [13], and details can be found in Appendix G. We show Lemmas 8 and 9 in Sections IX-A and IX-B, respectively.
@@ Appendix F Proof of Proposition 5: Error Analysis for the Protocol in Section VIII
First we can analyze the error of the rough estimate $\hat{\bm{p}}_{W}^{1}$. For each $w\in\mathcal{W}$, it is folklore that (cf. Theorem 4 in [20]),
$\mathbb{E}[|\hat{p}^{1}_{W}(w)-p_{W}(w)|^{p}|\mathcal{E}_{0}^{\complement}]=O\left(\left(\frac{p_{s}(w)(1-p_{s}(w))}{M_{0}}\right)^{\frac{p}{2}}\right)=O\left(\left(\frac{p_{W}(w)}{M_{0}}\right)^{\frac{p}{2}}\right).$ (41)
Then we derive the following preliminary result.
Lemma 15.
We have $\mathbb{P}\left[\frac{p_{W}(w)}{\hat{p}^{1}_{W}(w)}\geq 2\Big{|}\mathcal{E}_{0}^{\complement}\right]\leq O\left(\frac{1}{(M_{0}p_{W}(w))^{\frac{p}{2}}}\right)$.
Proof:
By the Markov inequality, we can obtain that
$\mathbb{P}\left[\frac{p_{W}(w)}{\hat{p}^{1}_{W}(w)}\geq 2\Big{|}\mathcal{E}_{0}^{\complement}\right]=\mathbb{P}\left[\frac{\hat{p}^{1}_{W}(w)}{p_{W}(w)}\leq\frac{1}{2}\Big{|}\mathcal{E}_{0}^{\complement}\right]$
$\leq$ $\mathbb{P}\left[\left|\hat{p}^{1}_{W}(w)-p_{W}(w)\right|\geq\frac{1}{2}p_{W}(w)\Big{|}\mathcal{E}_{0}^{\complement}\right]\leq\frac{2^{p}\mathbb{E}[\left|\hat{p}^{1}_{W}(w)-p_{W}(w)\right|^{p}|\mathcal{E}_{0}^{\complement}]}{p_{W}(w)^{p}}.$
Then by (41), we complete the proof. ∎
Next we return to the proof of Proposition 5.
Let $\mathcal{E}=\mathcal{E}_{0}\cup\left(\cup_{s=1}^{\lceil\log(\frac{k}{2^{l}})\rceil}\mathcal{E}_{s}\right)$. By the assumption that $m2^{l}\geq k^{2}$, we have $m^{\prime}2^{l}\geq k^{1.9}$. So we have
$\mathbb{P}[\mathcal{E}]\leq\sum_{s=0}^{\lceil\log(\frac{k}{2^{l}})\rceil}\mathbb{P}[\mathcal{E}_{s}]\leq e^{-\frac{m2^{l}}{64k}}+\sum_{s=1}^{\lceil\log(\frac{k}{2^{l}})\rceil}e^{-\frac{m^{\prime}2^{l}}{32(2^{l+s}+1)}}=O\left(e^{-\frac{m^{\prime}2^{l}}{128k}}\right)=O\left(\frac{k}{(m^{\prime}2^{l})^{\frac{p}{2}}}\right).$
Since $\|\hat{\bm{p}}^{2}_{W}-\bm{p}_{W}\|_{p}^{p}\leq 2$, we have
$\mathbb{E}[\mathds{1}_{\mathcal{E}}\cdot\|\hat{\bm{p}}^{2}_{W}-\bm{p}_{W}\|_{p}^{p}]\leq 2\mathbb{P}[\mathcal{E}]=O\left(\frac{k}{(m^{\prime}2^{l})^{\frac{p}{2}}}\right).$
It remains to show that for each $w\in\mathcal{W}$,
$\mathbb{E}\left[\mathds{1}_{\mathcal{E}^{\complement}}\cdot|p_{W}(w)-\hat{p}^{2}_{W}(w)|^{p}\right]=O\left(\frac{1}{(m^{\prime}2^{l})^{\frac{p}{2}}}+\left(\frac{p_{W}(w)}{m^{\prime}}\right)^{\frac{p}{2}}\right),$
then taking the summation over $w\in\mathcal{W}$ completes the proof. Similar to the proof of (41), for $w\in\mathcal{W}^{s}$ we have
$\mathbb{E}\left[|p_{W}(w)-\hat{p}^{2}_{W}(w)|^{p}\Big{|}\mathcal{E}^{\complement},w\in\mathcal{W}^{s}\right]=O\left(\mathbb{E}\left[\left(\frac{2^{l+s}p_{W}(w)}{m^{\prime}2^{l}}\right)^{\frac{p}{2}}\Big{|}\mathcal{E}^{\complement},w\in\mathcal{W}^{s}\right]\right)$
$=$ $O\left(\mathds{1}_{s>1}\cdot\mathbb{E}\left[\left(\frac{p_{W}(w)}{m^{\prime}2^{l}(\hat{p}^{1}_{W}(w)\vee\frac{1}{k})}\right)^{\frac{p}{2}}\Big{|}\mathcal{E}^{\complement},w\in\mathcal{W}^{s}\right]+\mathds{1}_{s=1}\cdot\mathbb{E}\left[\left(\frac{p_{W}(w)}{m^{\prime}}\right)^{\frac{p}{2}}\Big{|}\mathcal{E}^{\complement},w\in\mathcal{W}^{s}\right]\right).$
Taking the expectation, then we have
$\mathbb{E}\left[\mathds{1}_{\mathcal{E}^{\complement}}\cdot|p_{W}(w)-\hat{p}^{2}_{W}(w)|^{p}\right]=O\left(\mathbb{E}\left[\mathds{1}_{\mathcal{E}^{\complement}}\cdot\left(\frac{p_{W}(w)}{m^{\prime}2^{l}(\hat{p}^{1}_{W}(w)\vee\frac{1}{k})}\right)^{\frac{p}{2}}\right]+\left(\frac{p_{W}(w)}{m^{\prime}}\right)^{\frac{p}{2}}\right).$
To bound the first term, define the event $\mathcal{F}_{w}=\left\{\frac{p_{W}(w)}{\hat{p}^{1}_{W}(w)}\geq 2\right\}$. Then by Lemma 15 and $m2^{l}\geq k^{2}$, we have
$\mathbb{E}\left[\mathds{1}_{\mathcal{E}^{\complement}}\cdot\left(\frac{p_{W}(w)}{m^{\prime}2^{l}(\hat{p}^{1}_{W}(w)\vee\frac{1}{k})}\right)^{\frac{p}{2}}\right]$
$\leq$ $\mathbb{E}\left[\mathds{1}_{\mathcal{E}_{0}^{\complement}}\cdot\left(\frac{p_{W}(w)}{m^{\prime}2^{l}(\hat{p}^{1}_{W}(w)\vee\frac{1}{k})}\right)^{\frac{p}{2}}\right]$
$\leq$ $\mathbb{E}\left[\mathds{1}_{\mathcal{F}_{w}\cap\mathcal{E}_{0}^{\complement}}\left(\frac{p_{W}(w)}{m^{\prime}2^{l}\cdot\frac{1}{k}}\right)^{\frac{p}{2}}\right]+\mathbb{E}\left[\mathds{1}_{\mathcal{F}_{w}^{\complement}\cap\mathcal{E}_{0}^{\complement}}\left(\frac{p_{W}(w)}{m^{\prime}2^{l}\hat{p}^{1}_{W}(w)}\right)^{\frac{p}{2}}\right]$
$\leq$ $\mathbb{P}\left[\mathcal{F}_{w}\Big{|}\mathcal{E}_{0}^{\complement}\right]\cdot\left(\frac{kp_{W}(w)}{m^{\prime}2^{l}}\right)^{\frac{p}{2}}+O\left(\left(\frac{1}{m^{\prime}2^{l}}\right)^{\frac{p}{2}}\right)$
$\leq$ $O\left(\left(\frac{k}{m2^{l}p_{W}(w)}\right)^{\frac{p}{2}}\cdot\left(\frac{kp_{W}(w)}{m^{\prime}2^{l}}\right)^{\frac{p}{2}}\right)+O\left(\left(\frac{1}{m^{\prime}2^{l}}\right)^{\frac{p}{2}}\right)$
$=$ $O\left(\left(\frac{1}{m^{\prime}2^{l}}\right)^{\frac{p}{2}}\right).$
Hence we complete the proof of Proposition 5.

### v2

@@ Abstract
Consider the communication-constrained estimation of discrete distributions under $\ell^{p}$ losses, where each distributed terminal holds multiple independent samples and uses limited number of bits to describe the samples. We obtain the minimax optimal rates of the problem in most parameter regimes. An elbow effect of the optimal rates at $p=2$ is clearly identified. To show the optimal rates, we first design estimation protocols to achieve them. The key ingredient of these protocols is to introduce adaptive refinement mechanisms, which first generate rough estimate by partial information and then establish refined estimate in subsequent steps guided by the rough estimate. The protocols leverage successive refinement, sample compression, thresholding and random hashing methods to achieve the optimal rates in different parameter regimes. The optimality of the protocols is shown by deriving compatible minimax lower bounds.
@@ I Introduction
  • For $p>2$, we introduce additional sample compression methods to aid the adaptive refinement procedure. The methods compress the description for samples and reduce the communication budget, allowing more samples to be transmitted within limited budget. The resulting protocols can achieve the optimal rates for relatively large $n$. For $n=1$, an optimal non-interactive protocol can be designed, by exploiting random hash functions in the protocol. To show the optimality, we further establish a compatible lower bound that is strictly better than that in [15, 16],
@@ III-D Optimal Rates for $p>2$ and $n=1$
For $n=1$ and $p>2$, the lower bound can be derived by specializing Lemma 2, and the compatible upper bound is achieved by a non-interactive protocol, shown in the following thoerem.
@@
Let $p>2$ and $n=1$. We can design a non-interactive protocol that achieves the upper bound $R(m,1,k,l,p)\preceq\frac{k}{(m2^{l})^{\frac{p}{2}}}\vee\frac{1}{m^{\frac{p}{2}}}$. If $m(2^{l}\wedge k^{\frac{2}{p}})\geq k^{2}$, then $R(m,1,k,l,p)\succeq\frac{k}{(m2^{l})^{\frac{p}{2}}}\vee\frac{1}{m^{\frac{p}{2}}}$.
@@
The upper bound is by Proposition 5 in Section VIII. We sketch the proof here. A random hash function is used to compress the sample first, and then the estimate can be directly obtained by constructing and rescaling the histogram. See Section VIII for details of the protocol and its analysis. ∎
@@
Note that the central bound $\frac{1}{m^{\frac{p}{2}}}$ without the communication constraints is neglected by previous works [15, 16] (see Theorem 6 in [15] and Corollary 3.2 in [16]). Hence the lower bounds in both works are clearly not tight (for $p>2$). The work [16] further claimed that the lower bound $\frac{k}{(m2^{l})^{\frac{p}{2}}}\vee\frac{k^{1-\frac{p}{2}}}{m^{\frac{p}{2}}}$ is optimal (see Lemma 3.3 therein), but the sketch given there is too brief and not sufficient to describe a protocol that achieves the bound. In fact, given that the lower bound in [16] can be strictly improved, it is impossible to show its optimality. Moreover, constructing the protocol that achieves the optimal rates for $p>2$ is not that straightforward and needs additional ideas. We use random hashing to resolve the difficulty in this work.
@@ VIII The Protocol for $n=1$
In this section, we design a non-interactive protocol based on random hashing, which achieves the optimal rate for $n=1$. Similar to the discussion in Remark 5, it suffices to show the following proposition for $p\geq 2$.
@@
Let $p\geq 2$ and $n=1$. Then there exists a non-interactive protocol such that for any $\bm{p}_{W}\in\Delta_{\mathcal{W}}$, the protocol outputs an estimate $\hat{\bm{p}}_{W}$ satisfying $\mathbb{E}[\|\hat{\bm{p}}_{W}-\bm{p}_{W}\|_{p}^{p}]=O\left(\frac{k}{(m2^{l})^{\frac{p}{2}}}\vee\frac{1}{m^{\frac{p}{2}}}\right)$.
@@ VIII-A Motivation of the Protocol
The most natural idea is to first invoke the simulation protocol in [19] to output $M=O(\frac{m2^{l}}{k})$ samples from the distribution $\bm{p}_{\mathcal{W}}$ at the decoder side; then estimate $\bm{p}_{\mathcal{W}}$ using $M$ samples by a traditional central estimation method. It can achieve the optimal minimax rate $\frac{k}{m2^{l}}$ for $p=2$, and hence the optimal rate $\frac{k}{(m2^{l})^{\frac{p}{2}}}$ for $1\leq p\leq 2$. However, for $p\geq 2$, using $M$ i.i.d. samples to estimate the underlying distribution under the $\ell^{p}$ loss can only achieve a rate of $\frac{1}{M^{\frac{p}{2}}}=(\frac{k}{m2^{l}})^{\frac{p}{2}}$ , which leaves a gap with the lower bound $\frac{k}{(m2^{l})^{\frac{p}{2}}}$ by Lemma 1. The above naive protocol is not optimal and we can show that the lower bound $\frac{k}{(m2^{l})^{\frac{p}{2}}}$ is optimal.
@@
The aforementioned difficulty in estimation under $\ell^{p}$ losses can be overcome, by using a random hash function to compress the sample first, and then constructing and rescaling the histogram to obtain the estimate. No simulation step as in [19] is needed. Moreover, it is worth mentioning that the resulting protocol is non-interactive. The idea is similar to the second estimation stage in [9] for estimating a sparse distribution under communication constraints. Details of the protocol are presented as follows, and the error analysis can be found in Appendix F.
VIII-B The Non-interactive Protocol Based on Random Hashing for $n=1$
VIII-B1 Encoding
Let the $i$-th encoder generate a random hash function $h_{i}:\mathcal{W}\to\{0,1\}^{l}$, $i=1,...,m$ by shared randomness (i.e. $(h_{i}(w))_{w\in\mathcal{W}}$ are independent and $\mathbb{P}[h_{i}(w)=b]=2^{-l}$ for each $w\in\mathcal{W}$ and $b\in\{0,1\}^{l}$), so that the decoder can also generate $h_{i}$. Observing its sample $W_{i}$, the $i$-th encoder computes $B_{i}=h_{i}(W_{i})$ and sends it to the decoder.
VIII-B2 Decoding
Upon receiving $B_{i}$, the decoder then computes
$\hat{p}_{W}(w)=\frac{2^{l}}{2^{l}-1}\cdot\frac{\sum_{i=1}^{m}\mathds{1}_{h_{i}(w)=B_{i}}}{m}-\frac{1}{2^{l}-1}$ (15)
for each $w\in\mathcal{W}$ and outputs $\hat{\bm{p}}_{W}$.
@@ IX Lower Bounds
Lemma 6 is proved by exploiting the results for $p=1$ in [13], and details can be found in Appendix G. We show Lemmas 7 and 8 in Sections IX-A and IX-B, respectively.
@@ Appendix F Proof of Proposition 5: Error Analysis for the Protocol in Section VIII
We can analyze the error of the estimate $\hat{\bm{p}}_{W}$ as follows. Note that for each $w\in\mathcal{W}$ and $i=1,...,m$,
$\mathbb{P}[h_{i}(w)=B_{i}]=p_{W}(w)+\frac{1}{2^{l}}\left(1-p_{W}(w)\right).$
It is folklore that (cf. Theorem 4 in [20]),
$\mathbb{E}\left[\left|\frac{\sum_{i=1}^{m}\mathds{1}_{h_{i}(w)=B_{i}}}{m}-p_{W}(w)\right|^{p}\right]=O\left(\left(\frac{\mathbb{P}[h_{1}(w)=B_{1}]}{m}\right)^{\frac{p}{2}}\right)=O\left(\left(\frac{p_{W}(w)\vee\frac{1}{2^{l}}}{m}\right)^{\frac{p}{2}}\right).$
Then by (15), we have
$\mathbb{E}[|\hat{p}_{W}(w)-p_{W}(w)|^{p}]=O\left(\left(\frac{p_{W}(w)\vee\frac{1}{2^{l}}}{m}\right)^{\frac{p}{2}}\right)$
as well. By taking the summation over all $w\in\mathcal{W}$, we complete the proof of Proposition 5.